# Interpreting the Latent Structure of Operator Precedence in Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities but continue to struggle with arithmetic tasks. Prior works largely focus on outputs or prompting strategies, leaving the open question of the internal structure through which models do arithmetic computation. In this work, we investigate whether LLMs encode operator precedence in their internal representations via the open-source instruction-tuned LLaMA 3.2-3B model. We constructed a dataset of arithmetic expressions with three operands and two operators, varying the order and placement of parentheses. Using this dataset, we trace whether intermediate results appear in the residual stream of the instruction-tuned LLaMA 3.2-3B model. We apply interpretability techniques such as logit lens, linear classification probes, and UMAP geometric visualization. Our results show that intermediate computations are present in the residual stream, particularly after MLP blocks. We also find that the model linearly encodes precedence in each operator’s embeddings post attention layer. We introduce partial embedding swap, a technique that modifies operator precedence by exchanging high-impact embedding dimensions between operators.
@@ 1 Introduction
We employ a broad set of interpretability techniques including the logit lens, linear probes, partial embedding swaps, and geometric visualization via UMAP (nostalgebraist (2020); Alain & Bengio (2018); McInnes et al. (2020)). All experiments are conducted on the open-source, instruction-tuned LLaMA 3.2-3B model.
@@ 2 Related Works
Arithmetic Reasoning in Language Models. While LLMs demonstrate strong general reasoning abilities, numerous prior works have shown persistent inconsistencies for arithmetic tasks, especially when the prompted operations require multi-step computation and manipulation (Mirzadeh et al. (2024); Bubeck et al. (2023); Zhao et al. (2024)). These failures are particularly underlined in smaller models and remain poorly understood (Gangwar et al. (2025); Kim et al. (2024)). In a recent study (Boye & Moell (2025)), the authors evaluated arithmetic computations across multiple different models and observed frequent inconsistencies, such as over-reliance on numerical patterns and flawed logic, even when final answers were correct. In another study (Lewkowycz et al. (2022)), the authors have explored the use of prompting strategies, like chain-of-thought reasoning to support numerical computation.
Mechanistic Interpretability and Internal Components. Research in mechanistic interpretability has focused on identifying functional components, often referred to as “circuits”, that are responsible for specific model behaviors. A circuit refers to a subnetwork of model components that offers a faithful representation of how a model solves a particular task, such as a mathematical computation (Nainani et al. (2024)). For example, Zhang et al. (2024) investigates the internal structure of LLMs during arithmetic tasks, showing that fine-tuning a small subset of attention heads and MLPs can enhance arithmetic performance without compromising other abilities. These components are consistent across tasks and leads to overall better arithmetic performances. (Stolfo et al. (2023)) adopts a causal mediation framework to trace the influence of certain internal components that contribute the most to arithmetic prediction. However, their focus is limited in scope to focus on final predictions, rather than analyzing intermediate computational steps.
@@ 3.1 Model
Dataset Creation. We constructed a synthetic dataset consisting of arithmetic expressions with three operands and two binary operators. The dataset was designed to systematically test both syntactic and semantic precedence. We consider operands $a,b,c\in\{1,2,\dots,9\}$ and select operator pairs $(o_{1},o_{2})$ from the following mixed-precedence sets: { (+, *), (-, *), (+, /), (-, /) } Only mixed-precedence operator combinations are used to ensure meaningful precedence distinctions. For each combination of operands and operator pairs, we generate six structural variations: left-parenthesized: $(a\ o_{1}\ b)\ o_{2}\ c$, right-parenthesized: $a\ o_{1}\ (b\ o_{2}\ c)$, flipped left-parenthesized: $(a\ o_{2}\ b)\ o_{1}\ c$, flipped right-parenthesized: $a\ o_{2}\ (b\ o_{1}\ c)$, no-parentheses (natural order): $a\ o_{1}\ b\ o_{2}\ c$, no-parentheses (flipped): $a\ o_{2}\ b\ o_{1}\ c$. These expressions allow us to isolate the model’s handling of precedence both with and without explicit grouping via parentheses. For simplicity, prompts were selected such that all calculations, including intermediate steps, involved only positive whole numbers. In total 8547 prompts were created, but only prompts in which the model could predict the correct answer as the top logit were used to accurately examine model behavior. The model could answer 4401 equations correctly.
@@
Logit Lens to Trace Intermediate Computation. Before assessing how the model encodes operator precedence, we first examine whether it performs intermediate computations internally prior to generating the final output. For example, given the prompt “2 + 3 * 3 =”, we investigate whether the model computes the intermediate product 3 $\times$ 3 = 9 before arriving at the final result of 11. To account for linguistic variability in output tokens, we consider multiple surface forms (e.g., “11”, “eleven”, “eleventh”) as valid representations of the intermediate value. Using a dataset of 4,401 prompts for which the model produces correct final answers, we utilize the logit lens technique (nostalgebraist (2020)), projecting each layer’s residual stream through the model’s unembedding matrix to obtain logits over the vocabulary. We then extract the top 10 tokens by logit magnitude and check whether any of them match the expected intermediate result.
@@
Linear Probe for Latent Intermediate Computation. It is possible that the model performs intermediate computations internally, even when the corresponding value does not appear among the top logits. To further investigate this hypothesis, we train a linear probe (Alain & Bengio (2018)) to predict the intermediate value directly from the model’s activations at each layer, providing a complementary measure of whether such computations are linearly encoded.
@@
Operator Precedence in Embedding. To investigate whether operator precedence is encoded in specific dimensions of the model’s internal representations, we introduce partial embedding swap. Consider the prompt “3 + 4 * 5 = ”, where the correct evaluation yields the answer 23, adhering to standard arithmetic precedence. However, evaluating the expression strictly left-to-right results in the incorrect answer 35. We apply the algorithm described in Appendix A.2 to probe the embedding space. We selectively swap individual dimensions between the hidden representations of the ”+” and ”*” operator tokens. If such a perturbation induces the model to shift its prediction from 23 to 35, it provides evidence that those dimensions contribute to encoding precedence. This intervention is performed incrementally by swapping one dimension at a time and measuring the resulting change in the logit score assigned to the token “35”. Dimensions are then ranked according to their influence on increasing this logit. In the final phase of the experiment, we perform prefix swaps of the top k most influential dimensions to determine the minimal subset required to elevate “35” to the top-ranked prediction.
@@
Low-dimensional Projection of Embeddings. UMAP is a non-linear dimensionality reduction technique that facilitates the visualization and analysis of high-dimensional data by projecting it into a low-dimensional space (McInnes et al. (2020)). While specific embedding dimensions appeared to correlate with operator precedence, we employed UMAP to project the activation vectors corresponding to operator tokens from a curated set of prompts. These prompts varied only in the type and position of operators, allowing for controlled comparisons. Importantly, only prompts for which the model produced the correct answer were included, ensuring that the visualizations reflected meaningful internal representations. We labeled each operator using the following format: [position of the operator] [operator name] [precedence applied to the operator]. E.g., for a prompt  ”2 *  ( 3 + 4 ) = ”, the labels are ”1m2” and ”2p1”. The multiplication sign appears first in the expression but will be evaluated second, and the plus sign appears second in the equation but will be evaluated first.
@@
Logistic Probe to Identify Precedence. We employ linear probes (Alain & Bengio (2018)) to investigate whether operator precedence is linearly encoded in the internal representations of the model. We construct a dataset of arithmetic prompts containing two operations whose evaluation order determines the final result. To ensure alignment between model behavior and target labels, we include only correctly answered prompts. For each valid prompt, we extract activations at operator token positions from both pre- and post-attention within layer 0. A logistic regression classifier was trained to predict whether a given arithmetic operator (e.g., +, *) corresponds to the first or second operation to be evaluated in a two-step arithmetic expression (e.g., distinguishing between ”( 2 + 3 ) * 4 = ” and ”( 2 * 3 ) + 4 = ” ). Input to the probe consists of activation vectors associated with operator tokens in the prompt. The model is trained to predict either 0 or 1 based on whether the operator is evaluated first or second, respectively. For each probe, the activation vectors and binary labels are split into training and test sets (80%/20%). By comparing probe accuracy across the two extraction points (before and after attention), we assess whether and to what extent attention enhances the linear decodability of operator precedence.
@@ 4 Results and Analysis
Intermediate Calculation. Out of 4401 prompts, the intermediate calculation appeared 2799 times as the top logit, roughly 63.6%. The layers in which the calculation was discovered ranged from layer 16 - 27 for the LLaMA 3.2-3B model. This distribution is shown in Figure 1. To investigate whether it is the attention block or the MLP layer that introduces the intermediate logit, we apply the unembedding matrix to the outputs of each component at the layer where the intermediate logit first appears. In all cases, we find that the intermediate answer token’s logit becomes the top-ranked logit only after the MLP block, indicating that this component is responsible for producing the intermediate computation. Figure 1 strongly suggests that the intermediate calculation is linearly encoded in the model activations after layer 0.
Operator Precedence. We were able to successfully alter the model’s highest logit in multiple instances using partial embedding swap, examples shown in Appendix A.3. Our UMAP projection is shown in Figure 2. We found that operators who matched in both position and precedence were clustered near each other. Our logistic regression probe recieved 100% accuracy on our test set, strongly indicating the presence of operator precedence.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities but continue to struggle with arithmetic tasks. Prior works largely focus on outputs or prompting strategies, leaving the open question of the internal structure through which models do arithmetic computation. In this work, we investigate whether LLMs encode operator precedence in their internal representations via the open-source instruction-tuned LLaMA 3.2-3B model. We constructed a dataset of arithmetic expressions with three operands and two operators, varying the order and placement of parentheses. Using this dataset, we trace whether intermediate results appear in the residual stream of the instruction-tuned LLaMA 3.2-3B model. We apply interpretability techniques such as logit lens, linear classification probes, ablations, and UMAP geometric visualization. Our results show that intermediate computations are present in the residual stream, particularly after MLP blocks. We also find that the model linearly encodes precedence in each operator’s embeddings post attention layer. We introduce partial embedding swap, a technique that modifies operator precedence by exchanging high-impact embedding dimensions between operators.
@@ 1 Introduction
We employ a broad set of interpretability techniques including the logit lens, linear probes, ablations, partial embedding swaps, and geometric visualization via UMAP (nostalgebraist (2020); Alain & Bengio (2018); McInnes et al. (2020)). All experiments are conducted on the open-source, instruction-tuned LLaMA 3.2-3B model.
@@ 2 Related Works
2.1 Arithmetic Reasoning in Language Models
While LLMs demonstrate strong general reasoning abilities, numerous prior works have shown persistent inconsistencies for arithmetic tasks, especially when the prompted operations require multi-step computation and manipulation (Mirzadeh et al. (2024); Bubeck et al. (2023); Zhao et al. (2024)). These failures are particularly underlined in smaller models and remain poorly understood (Gangwar et al. (2025); Kim et al. (2024)). In a recent study (Boye & Moell (2025)), the authors evaluated arithmetic computations across multiple different models and observed frequent inconsistencies, such as over-reliance on numerical patterns and flawed logic, even when final answers were correct. In another study (Lewkowycz et al. (2022)), the authors have explored the use of prompting strategies, like chain-of-thought reasoning to support numerical computation.
2.2 Mechanistic Interpretability and Internal Components
Research in mechanistic interpretability has focused on identifying functional components, often referred to as “circuits”, that are responsible for specific model behaviors. A circuit refers to a subnetwork of model components that offers a faithful representation of how a model solves a particular task, such as a mathematical computation (Nainani et al. (2024)). For example, Zhang et al. (2024) investigates the internal structure of LLMs during arithmetic tasks, showing that fine-tuning a small subset of attention heads and MLPs can enhance arithmetic performance without compromising other abilities. These components are consistent across tasks and leads to overall better arithmetic performances. (Stolfo et al. (2023)) adopts a causal mediation framework to trace the influence of certain internal components that contribute the most to arithmetic prediction. However, their focus is limited in scope to focus on final predictions, rather than analyzing intermediate computational steps.
@@ 3.2 Dataset Creation
We constructed a synthetic dataset consisting of arithmetic expressions with three operands and two binary operators. The dataset was designed to systematically test both syntactic and semantic precedence. We consider operands $a,b,c\in\{1,2,\dots,9\}$ and select operator pairs $(o_{1},o_{2})$ from the following mixed-precedence sets: { (+, *), (-, *), (+, /), (-, /) } Only mixed-precedence operator combinations are used to ensure meaningful precedence distinctions. For each combination of operands and operator pairs, we generate six structural variations:
  1. 1.
    Left-parenthesized: $(a\ o_{1}\ b)\ o_{2}\ c$
  2. 2.
    Right-parenthesized: $a\ o_{1}\ (b\ o_{2}\ c)$
  3. 3.
    Flipped left-parenthesized: $(a\ o_{2}\ b)\ o_{1}\ c$
  4. 4.
    Flipped right-parenthesized: $a\ o_{2}\ (b\ o_{1}\ c)$
  5. 5.
    No-parentheses (natural order): $a\ o_{1}\ b\ o_{2}\ c$
  6. 6.
    No-parentheses (flipped): $a\ o_{2}\ b\ o_{1}\ c$
These expressions allow us to isolate the model’s handling of precedence both with and without explicit grouping via parentheses. For simplicity, prompts were selected such that all calculations, including intermediate steps, involved only positive whole numbers. In total 8547 prompts were created, but only prompts in which the model could predict the correct answer as the top logit were used to accurately examine model behavior. The model could answer 4401 equations correctly.
3.3 Logit Lens to Trace Intermediate Computation
@@
Before assessing how the model encodes operator precedence, we first examine whether it performs intermediate computations internally prior to generating the final output. For example, given the prompt “2 + 3 * 3 =”, we investigate whether the model computes the intermediate product 3 $\times$ 3 = 9 before arriving at the final result of 11. To account for linguistic variability in output tokens, we consider multiple surface forms (e.g., “11”, “eleven”, “eleventh”) as valid representations of the intermediate value. Using a dataset of 4,401 prompts for which the model produces correct final answers, we utilize the logit lens technique (nostalgebraist (2020)), projecting each layer’s residual stream through the model’s unembedding matrix to obtain logits over the vocabulary. We then extract the top 10 tokens by logit magnitude and check whether any of them match the expected intermediate result.
3.4 Linear Probe for Latent Intermediate Computation
@@
It is possible that the model performs intermediate computations internally, even when the corresponding value does not appear among the top logits. To further investigate this hypothesis, we train a linear probe (Alain & Bengio (2018)) to predict the intermediate value directly from the model’s activations at each layer, providing a complementary measure of whether such computations are linearly encoded.
3.5 Attention Layer Ablation to Probe Causal Contribution
To investigate whether intermediate computations are handled in the attention mechanisms of certain layers, we perform layer-wise attention ablations. For each expression, we selectively zero out the attention outputs at each layer, one layer at a time, while keeping all other components intact. For a given expression (e.g. “2 * (3 + 4) = “), we record the model’s output after ablating attention at the post-attention output before the MLP block. Ablation is implemented by setting the attention output tensor at a specific layer to zero before it is added to the residual steam. We run this procedure across the 27 transformer layers and evaluate the prediction accuracy. We then evaluate whether the intermediate result (e.g., 3 + 4 = 7) still appears in the residual stream via logit lens.
3.6 Operator Precedence in Embedding
@@
To investigate whether operator precedence is encoded in specific dimensions of the model’s internal representations, we introduce partial embedding swap. Consider the prompt “3 + 4 * 5 = ”, where the correct evaluation yields the answer 23, adhering to standard arithmetic precedence. However, evaluating the expression strictly left-to-right results in the incorrect answer 35. We apply the algorithm described in Appendix A.2 to probe the embedding space. We selectively swap individual dimensions between the hidden representations of the ”+” and ”*” operator tokens. If such a perturbation induces the model to shift its prediction from 23 to 35, it provides evidence that those dimensions contribute to encoding precedence. This intervention is performed incrementally by swapping one dimension at a time and measuring the resulting change in the logit score assigned to the token “35”. Dimensions are then ranked according to their influence on increasing this logit. In the final phase of the experiment, we perform prefix swaps of the top k most influential dimensions to determine the minimal subset required to elevate “35” to the top-ranked prediction. This method enables us to identify specific components of the embedding space that are linked to the model’s internal representation of operator precedence.
3.7 Low-dimensional Projection of Embeddings
@@
UMAP is a non-linear dimensionality reduction technique that facilitates the visualization and analysis of high-dimensional data by projecting it into a low-dimensional space (McInnes et al. (2020)). While specific embedding dimensions appeared to correlate with operator precedence, we employed UMAP to visualize the structure of the underlying representations. To this end, we projected the activation vectors corresponding to operator tokens from a curated set of prompts. These prompts varied only in the type and position of operators, allowing for controlled comparisons. Importantly, only prompts for which the model produced the correct answer were included, ensuring that the visualizations reflected meaningful internal representations. We labeled each operator using the following format: [position of the operator] [operator name] [precedence applied to the operator]. E.g., for a prompt  ”2 *  ( 3 + 4 ) = ”, the labels are ”1m2” and ”2p1”. The multiplication sign appears first in the expression but will be evaluated second, and the plus sign appears second in the equation but will be evaluated first.
3.8 Logistic Probe to Identify Precedence
@@
This work also investigates whether operator precedence is linearly encoded in the internal representations of the model. To this end, we employ linear probes (Alain & Bengio (2018)) trained on hidden state activations extracted before and after the first attention block (layer 0), enabling a comparative analysis of how the attention mechanism impacts the operator precendence. We construct a dataset of arithmetic prompts containing two operations whose evaluation order determines the final result. To ensure alignment between model behavior and target labels, we include only correctly answered prompts. For each valid prompt, we extract activations at operator token positions from both pre- and post-attention within layer 0. A logistic regression classifier was trained to predict whether a given arithmetic operator (e.g., +, *) corresponds to the first or second operation to be evaluated in a two-step arithmetic expression (e.g., distinguishing between ”( 2 + 3 ) * 4 = ” and ”( 2 * 3 ) + 4 = ” ). Input to the probe consists of activation vectors associated with operator tokens in the prompt. The model is trained to predict either 0 or 1 based on whether the operator is evaluated first or second, respectively. For each probe, the activation vectors and binary labels are split into training and test sets (80%/20%). By comparing probe accuracy across the two extraction points (before and after attention), we assess whether and to what extent attention enhances the linear decodability of operator precedence.
@@
Figure 2: Shows impact of attention ablation on output accuracy across 19 layers of the model. Early layers, particularly layer 0, showed the most pronounced effect on accuracy, suggesting a key role in establishing the structural interpretation of arithmetic expressions.
Figure 3: Shows impact of attention ablation on intermediate computation observed through Logit Lens across 4401 equations. Layers 9, 13, and 19 had the most decreases in observed intermediate values, suggesting a key role in establishing intermediate computations.
@@ 4 Results and Analysis
4.1 Intermediate Calculation
Out of 4401 prompts, the intermediate calculation appeared 2799 times as the top logit, roughly 63.6%. The layers in which the calculation was discovered ranged from layer 16 - 27 for the LLaMA 3.2-3B model. This distribution is shown in Figure 1. To investigate whether it is the attention block or the MLP layer that introduces the intermediate logit, we apply the unembedding matrix to the outputs of each component at the layer where the intermediate logit first appears. In all cases, we find that the intermediate answer token’s logit becomes the top-ranked logit only after the MLP block, indicating that this component is responsible for producing the intermediate computation. Figure 1 strongly suggests that the intermediate calculation is linearly encoded in the model activations after layer 0. Notably, ablating attention in early layers significantly impaired performance: ablating layer 0 led to incorrect answers in 95.7% of cases, with layers 1, 5, and 7 also showing elevated error rates of 42.2%, 21.1%, and 24.2%, respectively (Figure 2). Layers after 19 showed limited effects. Across 19 layers, attention ablation significantly reduced the occurrence of intermediate computations observed through the logit lens, with the most pronounced effects in layers 5, 11, and 19. These findings indicate that the model relies on the attention mechanism to construct intermediate arithmetic representations throughout its depth, with layers 5, 11, and 19 playing a particularly critical role. (Figure 3) .
4.2 Operator Precedence
We were able to successfully alter the model’s highest logit in multiple instances using partial embedding swap, examples shown in Figure 5. Our UMAP projection is shown in Figure 4. We found that operators who matched in both position and precedence were clustered near each other. Our logistic regression probe recieved 100% accuracy on our test set, strongly indicating the presence of operator precedence.